please show correct answer, thank you
Find and sketch the domain of the function. f(x, y) = √y+√4x² - y2 у 0 0 4 X 0 2 0 Type here to search a i
Answer #1

Ques Given fla,y)= sy + 2 4 We will find Domain of f(x,y) Domain Value of azy such that function frony) is defined is definedyzo and 4-12-y? > 0 =) x² + y² = 4 Ganotion a is Unit Disk of badius 2. Sketch - I a x² + y2 < (2) 2 y x²7y² =4 - Sketch I T(Take Intehsection of Sketch - I 2 we get كه I (-2,0) is Correct option